Which is a zero of the function f(x)=x^2+6x-7?
Answer by nivine(1)

x^2-6x-7=0
when he ask this question we equal the equation to zer0
then we try to find the roots for the equation
the roots means we solve the equation in this shape
( )( )=zero
so
(x-7)(x+1)=0 to make sure u can multiply the two pranticese u will get the same equation
then each brantic. = 0 so x-7=0 so x=-7
and x+1=0 so x=-1